Alpha is planning on merging with Beta. Alpha will pay Beta's shareholders the current value of their equity in shares of Alpha. Alpha currently has 4,200 shares of equity outstanding at a market price of £40 a share. Beta has 2,500 shares outstanding at a price of £18 a share. The after-merger earnings will be £8,800. What will the earnings per share be after the merger?
Answer #1

Calculation of number of shares after merger:

Merger consideration = No. of shares of beta*price per beta's share = 2,500*18= GBP 45,000

No. of shares of alpha to be issued to beta's shareholders= merger consideration/price per alpha's share= 45,000/40= 1,125 shares

Total no. of shares outstanding post-merger= Initial shares of alpha + new shares of alpha issued to beta's shareholders= 4,200+1,125= 5,325 shares

Earnings per share post-merger= After-merger earnings/total no. of post-merger shares = 8,800/5,325= GBP 1.65 per share